# UtilityMax Prompting: A Formal Framework for Multi-Objective Large Language Model Optimization

### 12 Mar 2026 v1

@@ 1 Introduction
The rest of this paper is organized as follows: in Section 2 we cover the general methodology behind the UtilityMax framework; in Section 3 we extend the framework to handle dependencies between binary chance nodes; in Section 4 we provide a prompting template under UtilityMax that practitioners can use; in Section 5 we validate UtilityMax on a multi-objective movie recommendation task using the MovieLens 1M dataset Harper and Konstan (2015), demonstrating consistent improvements in precision and Normalized Discounted Cumulative Gain (NDCG) over natural language baselines across three frontier models (Claude Sonnet 4.6, GPT-5.4, and Gemini 2.5 Pro); and we conclude with final remarks and directions for future research in Section 6.
@@ 2 Framework
Let $K$ represent the LLM’s knowledge. This includes all internal knowledge stored through its parameters as well as external knowledge such as context through chat history or online research. Given $K$, we define the following influence diagram:
Let $A\mid K$ be a decision node that represents the space of all possible LLM’s answers given $K$. An answer $a\in A$ is decided by the LLM. Let $\{X_{1},X_{2},X_{3},...,X_{n}\}$ be a set of chance nodes with $n\geq 1$. We require:
@@
    The dependency structure over $\{A,X_{1},...,X_{n}\}$ forms a directed acyclic graph (DAG) with $A$ as the singular root node.
@@
Suppose further that we define a multiplicative utility function $U(X_{1},...,X_{n})=\prod_{i=1}^{n}f_{i}(X_{i})$. Then the expected utility factorizes as:
@@
This setup is shown visually in Figure 1.
@@
Figure 1: Influence diagram for UtilityMax Prompting showing decision node AA, chance nodes $\{X_{1},X_{2},X_{3},...,X_{n}\}$, and utility node $U$.
Then the task of the LLM is to find $a^{*}\in A$ that maximises $\mathbb{E}[U\mid A=a^{*}]$.
3 Binary Chance Nodes
In the special case where the chance nodes $\{X_{1},X_{2},\ldots,X_{n}\}$ are all binary random variables, we can relax the conditional independence assumption. Let $pa(X_{i})$ denote the parents of $X_{i}$ in the DAG, and suppose that if any $X_{j}=0$ for $X_{j}\in pa(X_{i})$ then $P(X_{i}=1\mid pa(X_{i}),A)=0$. That is, each node is deterministically gated by its parents so that a child node can only be active if all its parents are active.
Denote by $pa(X_{i})=1$ the event that $X_{j}=1$ for all $X_{j}\in pa(X_{i})$. Under the gating assumption, all terms in the expectation where any parent takes the value zero vanish, and the expected utility reduces to:
$\mathbb{E}[U\mid A]=\prod_{i=1}^{n}P(X_{i}=1\mid pa(X_{i})=1,A)$ (2)
where we define $f_{i}(X_{i})=X_{i}$ so that $\mathbb{E}[X_{i}\mid pa(X_{i})=1,A]=P(X_{i}=1\mid pa(X_{i})=1,A)$. Each term is estimated independently by the LLM, preserving tractability without requiring conditional independence.
4 Prompting Template
@@
5 Experimental Setup
To validate our prompting framework we run experiments on the MovieLens 1M dataset Harper and Konstan (2015). Given the first $100$ movies that a user has rated between $1$ and $5$, the LLM’s task is to recommend the top $10$ movies for that user from the next $50$ in their watched list.
@@ 5 Experimental Setup
    UtilityMax: A prompt is constructed based on the template of Section 4 with three random variables - a categorical random variable $S$ for the predicted score, a binary random variable $G_{1}$ for the comedy genre, and a binary random variable $G_{2}$ for the romance genre. Therefore, the objective is:
@@ 6 Final Remarks
The effectiveness of UtilityMax depends on two key factors. First, the underlying model must be expressive enough to produce well-calibrated probability estimates for the objective components. Our experiments suggest that current frontier models meet this requirement, at least on the MovieLens recommendation task.
@@
Several directions for future research present themselves. Perhaps the most practically important is automating the construction of the UtilityMax prompt. Specifically, developing a method by which an LLM can extract the relevant variables from a natural language task description and format them into the UtilityMax framework automatically. This would remove the requirement for a human designer to manually specify the optimization objective.
A second important direction is further relaxing the conditional independence assumption on the chance nodes given $A$. While this assumption yields a tractable objective and is sufficient to capture a wide range of tasks, it precludes dependencies between chance nodes that may be necessary in certain domains. Extending the framework beyond the gating mechanism introduced in Section 3 to handle such dependencies while preserving tractability remains an open problem.

### 26 Mar 2026 v2

@@ 1 Introduction
The rest of this paper is organized as follows: in Section 2 we cover the general methodology behind the UtilityMax framework; in Section 3 we provide guidelines as well as a prompting template under UtilityMax for practitioners; in Section 4 we validate UtilityMax on a multi-objective movie recommendation task using the MovieLens 1M dataset Harper and Konstan (2015), demonstrating consistent improvements in precision and Normalized Discounted Cumulative Gain (NDCG) over natural language baselines across three frontier models (Claude Sonnet 4.6, GPT-5.4, and Gemini 2.5 Pro); and we conclude with final remarks and directions for future research in Section 5.
@@ 2 Framework
We formalise UtilityMax as an influence diagram in which the LLM’s answer is the sole decision variable. The tractability of the expected utility depends on the dependency structure among the chance nodes. We describe two sufficient conditions below.
2.1 Conditional Independence
In the simplest case, we assume the chance nodes are mutually conditionally independent given the LLM’s answer. Formally, let $K$ represent the LLM’s knowledge. This includes all internal knowledge stored through its parameters as well as external knowledge such as context through chat history or online research. Given $K$, we define the following influence diagram:
Let $A\mid K$ be a decision node that represents the space of all possible LLM’s answers given $K$. An answer $a\in A$ is decided by the LLM. Let $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{n}\}$ be a set of chance nodes with $n\geq 1$. We require:
@@ 2.1 Conditional Independence
    The dependency structure over $\{A\}\cup\mathbf{X}$ forms a directed acyclic graph (DAG) with $A$ as the singular root node.
@@
Suppose further that we define a multiplicative utility function $U(\mathbf{X})=\prod_{i=1}^{n}f_{i}(X_{i})$. Then the expected utility factorizes as:
@@
This setup is shown visually in Figure 1. Then the task of the LLM is to find $a^{*}\in A$ such that $\mathbb{E}[U\mid A=a^{*}]$ is maximised.
@@
Figure 1: Influence diagram for UtilityMax Prompting showing decision node AA, chance nodes $\{X_{1},X_{2},\ldots,X_{n}\}$, and utility node $U$.
2.2 Binary Gating
If all internal (non-leaf) chance nodes are binary random variables, we can relax the conditional independence assumption. Let $pa(X_{i})$ denote the parents of $X_{i}$ in the DAG, and suppose that if any $X_{j}=0$ for $X_{j}\in pa(X_{i})$ then $P(X_{i}=1\mid pa(X_{i}),A)=0$. That is, each internal node is deterministically gated by its parents so that a child node can only be active if all its parents are active. Leaf nodes are not required to be binary and may take continuous or categorical values.
Denote by $pa(X_{i})=1$ the event that $X_{j}=1$ for all $X_{j}\in pa(X_{i})$. Under the gating assumption, all terms in the expectation where any parent takes the value zero vanish. Let $\mathbf{L}\subset\mathbf{X}$ denote the set of leaf nodes and $\mathbf{I}=\mathbf{X}\setminus\mathbf{L}$ denote the set of internal nodes. We further assume that for $X_{i}\in\mathbf{I}$, $f_{i}(X_{i})=X_{i}$. This is without loss of generality since we require $f_{i}(0)=0$ for tractability. Then the expected utility reduces to:
$\mathbb{E}[U\mid A]$ $=\prod_{i=1}^{n}\mathbb{E}[f_{i}(X_{i})\mid pa(X_{i})=1,A]$
$=\prod_{i:\,X_{i}\in\mathbf{I}}P(X_{i}=1\mid pa(X_{i})=1,A)\times\prod_{j:\,X_{j}\in\mathbf{L}}\mathbb{E}[f_{j}(X_{j})\mid pa(X_{j})=1,A].$ (2)
The first product captures the binary gating probabilities across internal nodes. The second captures the expected utility contribution of each leaf node, conditioned on all its parents being active. Each term is estimated independently by the LLM, preserving tractability without requiring conditional independence. In the special case where all nodes are binary and $f_{i}(X_{i})=X_{i}$ for all $i\in\{1,\ldots,n\}$, the expected utility reduces to:
$\mathbb{E}[U\mid A]=\prod_{i=1}^{n}P(X_{i}=1\mid pa(X_{i})=1,A).$ (3)
3 Prompting
3.1 Guidelines
The effectiveness of UtilityMax depends not only on the structure of the influence diagram but also on the precision of the variable definitions. There is an inherent trade-off: a precisely defined variable constrains the LLM to a narrow target, reducing ambiguity in its estimates but potentially missing relevant signal that falls outside the strict definition.
A loosely defined variable gives the LLM more latitude to draw on its broader knowledge, at the cost of reintroducing the same ambiguity that UtilityMax is designed to eliminate. In general, a variable definition is considered more precise the less the LLM needs to rely on natural language interpretation to estimate its probability.
The right balance depends on the task. When the outcome can be well-specified, a more precise definition is generally preferred. When the concept is inherently subjective, or the designer wants to leverage the LLM’s broader judgment, some looseness may be acceptable.
To maximise the accuracy between the real-world targets and the LLM’s estimates, variables should be defined to describe an observable event whose outcome can be measured. For example, consider a binary variable where success is defined as “collaboration occurs,” compared to “at least one person contacts me within two weeks about running a joint experiment.” The latter is more concrete, so the LLM can form a more meaningful probability estimate.
3.2 Template
@@
4 Experiments
To validate our prompting framework, we run experiments on the MovieLens 1M dataset Harper and Konstan (2015). Given the first $100$ movies that a user has rated between $1$ and $5$, the LLM’s task is to recommend the top $10$ movies for that user from the next $50$ in their watched list.
@@ 4 Experiments
    UtilityMax: A prompt is constructed based on the template of Section 3.2 with three random variables - a categorical random variable $S$ for the predicted score, a binary random variable $G_{1}$ for the comedy genre, and a binary random variable $G_{2}$ for the romance genre. Therefore, the objective is:
@@ 5 Final Remarks
The effectiveness of UtilityMax depends on two key factors. First, the underlying model must be expressive enough to produce well-calibrated probability estimates for the objective components. This requires $K$ to be sufficiently rich – drawing on the model’s internal knowledge, externally provided context, or iterative dialogue. While UtilityMax requires no task exemplars and is therefore zero-shot in that sense, enriching $K$ through additional context or chat iteration can improve performance without altering the formal objective structure.
@@
Several directions for future research present themselves:
  1. 1.
    Perhaps the most practically important is automating the construction of the UtilityMax prompt. Specifically, developing a method by which an LLM can extract the relevant variables from a natural language task description and format them into the UtilityMax framework automatically. This would remove the requirement for a human designer to manually specify the optimization objective.
  2. 2.
    A second direction is identifying further DAG structures beyond the binary gating mechanism of Section 2.2 that yield tractable expected utility functions.
  3. 3.
    In the current framework, the LLM relies solely on $K$ to estimate each component probability from scratch. However, in many real-world settings historical data may be available that could be used to produce empirically calibrated prior estimates. Machine learning methods for parameter estimation in probabilistic graphical models could be applied to learn initial probability estimates for each node in the DAG. These data-driven priors would then serve as a starting point that the LLM refines using its broader knowledge and contextual understanding given its answer $A$. This is analogous to Bayesian updating where the data provides the prior and the LLM provides the likelihood update.
  4. 4.